# Tactile Displays Driven by Projected Light

### v1

@@ Introduction
We next investigated the dynamic response of the display under rapid cyclic photostimulation. Rapid stimulation produces a multimodal displacement response $d(t)=\delta(t)+\delta_{o}(t)$ with slowly varying and oscillating components, $\delta(t)$ and $\delta_{o}(t)$ respectively (Fig. 3C). The slow component, $\delta(t)$, is due to the retention of heat in the cavity across cycles, as occurs when the inter-pulse gap time, $t_{g}$, is shorter than $\tau.$ This component attenuates in amplitude as $t_{g}$ increases; in steady-state, $\delta_{o}/\delta_{1}=\tau/t_{g}$ (Fig. 3D, red). The oscillating component, which is felt as a vibration, exhibits an opposite trend, described by $\delta_{pp}/\delta_{1}=1-\exp(-t_{g}/\tau)$ (Fig. 3D, blue). Across a range of values of $t_{g}$, both components fall within the perceptually salient amplitude range  (?). Thus, by adjusting the pulse gap time, $t_{g}$, we can control the relative proportion of slow and vibratory tactile feedback.
@@
The pulse rate of cyclic photostimulation $f=1/T$ is the inverse of the pulse period, $T=t_{p}+t_{g}$. With optical power $P_{L}$ and duty cycle $\nu=t_{p}/T$ held constant, the quasi-static displacement, $\delta_{o},$ is nearly invariant with respect to variations in pulse rate $f$ (Fig. S10). This occurs because the average absorbed power $\overline{P}=\nu\varepsilon P_{L}$ is determined by the duty cycle and optical power. In contrast, because $t_{p}$ and $t_{g}$ both decrease as $f$ increases, the oscillation amplitude, $\delta_{pp},$ decreases (Fig. 3E). Nonetheless, even rapid pulsing at $f=200$ Hz yields an oscillation amplitude of $\delta_{pp}=8.4$ $\mu$m, which is large enough to be perceived via touch. For comparison, this amplitude is about 25 dB higher than the perceptual detection threshold for 200 Hz vibrations at the fingertip  (?).
@@
While the response rate of cyclically driven pixels is constrained by the thermal response time scale, our display can also render tactile patterns by rapidly scanning the light source, at 100% duty cycle, across arrays of pixels in sequence. In sequential scanning, the maximum display rate $N$, in pixels per second, is dictated by the pulse time, $t_{p},$ needed to achieve a specified displacement amplitude, $d$, at the operating power level, $P_{L}$. The maximum pixel rate $N$ thus decreases with increasing displacement (Fig. 3F). At power $P_{L}=2.5$ W, specifying displacement to be $d=$ 50 $\mu$m results in a maximum display rate of $N=$ 217 pixels/s, while for $d=$ 400 $\mu$m the maximum rate is $N=26.5$ pixels/s.
@@
Figure 4: Fig. 4. Perception of Dynamic Tactile Pattern Display. A) Experiment 1: Perception of linear tactile motion patterns. All participants correctly reported the motion direction in 100% of trials. B) Experiment 2: Perception of rotational motion for motion speeds $v=$ 16, 32, and 64 mm/s. Mean response accuracy: 94.7%. Black dots show the overall mean, and grey dots show participant mean. C) Experiment 3: Spatial localization of actuated pixels near the finger. Mean accuracy: 78.4% (99% within 1 pixel distance). Mean localization error: 0.17 mm. D) Experiment 4: The perceived intensity of pulse train feedback from a single pixel increased linearly with optical power $P_{L}$ ($\alpha=0.022$, $\beta=-0.42$, $r^{2}=0.98$). Stimulus parameters: $N=20$ pulses, $t_{p}=t_{g}=$ 25 ms, $2.50.5<P_{L}<$ 2.5 W. E) Experiment 5: Discrimination of temporal patterns. ”Odd-One-Out” task to identify which of three sensations was different. Total accuracy: 89.6%. F) Experiment 6: Perception of multi-point tactile patterns felt at any two of four fingers. Participants reported the locations that were stimulated with 96.3% accuracy.
This tactile display employs a physical operating principle that diverges from those used in prior haptic technologies. The forces and displacements that can be produced by individual pixels (up to 55 mN and 0.97 mm) are well within the perceivable range  (?). These basic capabilities led us to investigate the potential of using multiple pixels to render a variety of perceptually distinct tactile patterns. We conducted six experiments assessing the perception of linear and rotational motion patterns, tactile localization, the scaling of perceived intensity with power, the perception of temporal patterns, and multi-finger tactile perception (Figs. 4A-F). The same ten participants completed all six experiments without training. In post-hoc surveys, no participants reported perceiving thermal sensations or seeing any optical stimuli.
@@ Conclusion
Our tactile display technique exploits the advantages of projected light for power transmission, addressing, and actuation. Light is thermodynamically transduced into forces and displacements via energetically passive surfaces populated with optotactile pixels. The actuation mechanism, layered architecture, and simplicity of fabrication make it straightforward to produce displays with dozens to thousands of pixels, with similar investments in material cost and fabrication time. We present implementations with up to 1,511 pixels – significantly more than any previous device capable of meeting similar performance requirements. Our technique furthermore yields tactile patterns that can be not only felt, but also seen, enabling multisensory tactile displays. These displays hold promise for applications ranging from automotive interfaces to virtual reality controllers. In order for them to achieve widespread adoption, higher pixel display rates are needed. This could be achieved using optimized architectures that exploit fundamental knowledge about thermodynamic device efficiency, and by leveraging emerging high-power optical projection technologies for additive manufacturing and automotive applications  (?, ?).
@@ Acknowledgments
We gratefully acknowledge Vijay Kumar, Yangying Zhu, Shantal Adajian, Bolin Liao, and Alban Sauret for assistance with characterization experiments. We thank Neeli Tummala for valuable comments. We acknowledge the Innovation Workshop and Microfluidics Lab in the UCSB California NanoSystems Institute, the UCSB MRSEC facility (NSF DMR 0520415), and the UCSB Nanofabrication facility, which were used in this research.
Author Contributions: M.L. and Y.V., designed research; M.L, G.R., D.G. and Y.V. performed research; M.L., E.H., and Y.V. analyzed data; M.L, and Y.V. wrote the paper.

### v3

@@ Introduction
We next investigated temporal parameters governing the dynamic response of the display under cyclic photostimulation. Rapid stimulation produces a multimodal displacement response $z(t)=d(t)+\delta(t),$ with slowly varying and oscillating components, $d(t)$ and $\delta(t)$ respectively (Fig. 3C). The slow component, $d(t)$, is due to the retention of heat in the cavity across cycles, as occurs when the inter-pulse gap time, $t_{g}$, is shorter than $\tau.$ The amplitude of $d(t)$ decreases as $t_{g}$ increases; in steady-state, $d/d_{1}=\tau/t_{g}$ (Fig. 3D, red), where $d_{1}$ is the first pulse amplitude. The amplitude of the oscillating component, $\delta_{pp}$, exhibits an opposite trend, described by $\delta_{pp}/\delta_{1}=1-\exp(-t_{g}/\tau)$ (Fig. 3D, blue). Across a range of values of $t_{g}$, both components fall within the perceptually significant amplitude range  (?). Thus, by adjusting the pulse gap time, $t_{g}$, we can control the relative proportion of slow and vibratory tactile feedback.
@@
Altering the pulse rate, $f=1/(t_{p}+t_{g}),$ differentially modulates the amplitude of the two components of displacement. At constant power and duty cycle, $\nu=t_{p}/T,$ the amplitude of the slow component, $d$, is nearly invariant with respect to variations in pulse rate $f$ (Fig. S10). This occurs because the average absorbed power $\overline{P}=\nu\varepsilon P_{L}$ is determined by the duty cycle and optical power. In contrast, because $t_{p}$ and $t_{g}$ both decrease as $f$ increases, the oscillation amplitude, $\delta_{pp},$ decreases (Fig. 3E). Nonetheless, even rapid pulsing at $f=200$ Hz yields an oscillation amplitude of $\delta_{pp}=8.4$ $\mu$m, which is sufficient to be perceived via touch. [^3] For comparison, this amplitude is about 25 dB higher than the perceptual detection threshold for 200 Hz vibrations at the fingertip  (?).
@@
While the response rate of cyclically driven pixels is constrained by the thermal response time scale, our display can also render tactile patterns by rapidly scanning the light source across arrays of pixels in sequence. In sequential scanning, the maximum display rate $N$, in pixels per second, is dictated by the pulse time, $t_{p},$ needed to achieve a specified amplitude, $\delta_{pp}$, at the operating power level, $P_{L}$. The maximum pixel rate $N$ thus decreases with increasing displacement (Fig. 3F). At power $P_{L}=2.5$ W, specifying displacement to be $d=$ 50 $\mu$m results in a maximum display rate of $N=$ 217 pixels/s, while for $d=$ 400 $\mu$m the maximum rate is $N=26.5$ pixels/s. Higher rates can be achieved by increasing optical power or employing multiple optical sources.
@@
Figure 4: Fig. 4. Perception of Dynamic Tactile Pattern Display. (A) Experiment 1: Perception of linear tactile motion patterns. All participants correctly reported the motion direction in 100% of trials. (B) Experiment 2: Perception of rotational motion for motion speeds $v=$ 16, 32, and 64 mm/s. Mean response accuracy: 94.7%. Black dots show the overall mean, and grey dots show participant mean. (C) Experiment 3: Spatial localization of actuated pixels near the finger. Mean accuracy: 78.4% (99% within 1 pixel distance). Mean localization error: 0.17 mm. (D) Experiment 4: The perceived intensity of pulse train feedback from a single pixel increased linearly with optical power $P_{L}$ ($\alpha=0.022$, $\beta=-0.42$, $r^{2}=0.98$). Stimulus parameters: $N=20$ pulses, $t_{p}=t_{g}=$ 25 ms, $2.50.5<P_{L}<$ 2.5 W. (E) Experiment 5: Discrimination of temporal patterns. ”Odd-One-Out” task to identify which of three sensations was different. Total accuracy: 89.6%. (F) Experiment 6: Perception of multi-point tactile patterns felt at any two of four fingers. Participants reported the locations that were stimulated with 96.3% accuracy.
This tactile display employs a physical operating principle that diverges from those used in prior haptic technologies. The forces and displacements that can be produced by individual pixels (up to 55 mN and 0.97 mm) are well within the perceivable range  (?). These basic capabilities led us to investigate the potential of using multiple pixels to render a variety of perceptually distinct tactile patterns. We conducted six experiments assessing the perception of linear and rotational motion patterns, tactile localization, the scaling of perceived intensity with power, the perception of temporal patterns, and multi-finger tactile perception (Figs. 4A-F). The same ten participants completed all six experiments without training.
@@
In post-hoc surveys, no participants reported perceiving thermal sensations or seeing any optical stimuli.
@@ Conclusion
Our tactile display technique exploits the advantages of projected light for power transmission, addressing, and actuation. Light is directly transduced into forces and displacements via energetically passive surfaces populated with optotactile pixels. The actuation mechanism, layered architecture, and simplicity of fabrication make it straightforward to produce displays with dozens to thousands of pixels, with similar investments in material cost and fabrication time. We present implementations with up to 1,511 pixels – significantly more than any previous device capable of similar performance, notably in response time and amplitude. Our technique furthermore yields tactile patterns that can be not only felt, but also seen, enabling multisensory tactile displays. These displays hold promise for applications ranging from automotive interfaces to virtual reality controllers. In order for them to achieve widespread adoption, higher pixel display rates are needed. This could be achieved using optimized architectures that exploit fundamental knowledge about thermodynamic device efficiency, and by leveraging high-power optical projection technologies like those being developed for additive manufacturing and automotive applications  (?). These improvements could enable widespread deployment of high-definition projected light tactile displays that lend physical form to otherwise intangible digital data.
Materials and methods
Display fabrication
The same manufacturing steps were used for displays with 1 to 1,511 pixels. The display surface is a layered assembly. In order from base to top layer, the materials are: Transparent acrylic sheet, one-side adhesive polysiloxane (PS) sheet (thickness 0.5 mm), thick pyrolytic graphite sheet (thickness 17 $\mu$m), one-side adhesive PS sheet, cast silicone elastomer (thickness 0.25 mm). Double-sided adhesive tape (2477, 3M, USA) was used to affix the lower PS sheet to the acrylic. The thickness of the acrylic may be selected based on application requirements. For the flexible display, the thickness was 0.01”, and for the rigid display, it was 1/16”. Supplier information, and representative costs for the materials, are reported in Table S5.
The PS sheets were patterned via laser cutting to form a grid of cylindrical cavities ($r_{\mathrm{cavity}}$ = 1.5 mm, pitch = 4 mm; Fig. S5). The PGS was patterned using a vinyl cutter to form a grid of rectangular photoabsorbers (2 mm x 1.5 mm) aligned with the cavities (Fig. S5). The acrylic was laser cut to form the square geometry; for the 1,511 display, the dimensions are 162 mm x 162 mm (Fig. S5). The PGS adhesive was dissolved via acetone processing, and aligned between PS layers. For the cast elastomer, the two silicone elastomer component reagents were mixed, degassed, and placed into a mold with depth 250 $\mu$m . It was subsequently cured at room temperature at room temperature for 24 hours. Because the resulting elastomer sheet is tacky, no adhesive was needed to affix it to the PS layer.
Photoabsorber Thickness and Density Characterization
The thickness of the PGS photoabsorber was measured via profilometry and confocal microscopy. For the profilometer measurement (DektakXT, Bruker), the adhesive side of the PGS was applied to a glass slide. Measurements from four distinct samples of the combined thickness of PGS and adhesive were taken, yielding $h_{PGS,adh}=23.2\pm 1.2\mu$m (Fig. S16). We characterized the adhesive thickness using confocal microscopy (DSX1000, Olympus), obtaining $h_{adh}=6.5\pm 1\mu$m. Subtracting the two thicknesses, yielded the PGS photoabsorber thickness, $h_{PGS}=16.7\pm 2.2\mu$m. The result was 6.7 $\mu$m larger than the thickness of 10 $\mu$m specified by the manufacturer. The PGS surface roughness was ±$\pm 1.5$ $\mu$m. The mass density of the PGS material, after adhesive removal, was determined using a microbalance (XPR105, Mettler Toledo), and dimensional measurements of surface area ($25\pm 5$ mm2) and thickness $h_{PGS}$ (as noted above). This yielded a PGS density value of $\rho_{PGS}=1962\pm 650$ kg/m3, which is consistent with the manufacturer specified value of 2130 kg/m3.
Optical stage
The optical stage included a 2-axis galvo mirror system (Laserwinkel, Netherlands), a 4W 450 nm diode laser (LT-20W-A, Laser Tree), a collimating lens (Qiaoba), and an F-theta lens (Opex). The optical components were arranged in a 3D printed fixture. The laser path proceeded through the collimating lens, galvo, F-theta lens, and through the optical window of the addressed pixel, onto the respective photoabsorber (Fig. S6). The optical stage was configured in an optically sealed enclosure with an aperture located at the top, where the display surface is positioned. Representative costs of components used to construct the optical stage are given in Table S6.
The laser was powered with 12 V using a laboratory supply. Optical power output was modulated by PWM signal (PWM frequency: 5 kHz) supplied to the laser by a data acquisition unit (NI SCB-68A) under computer control.
The maximum instantaneous optical power incident on the sample was 2.45 W, as measured via bolometer (S212A, Thorlabs, USA). This measured power accounted for transmission losses through the mirrors and lenses. Transmission through the acrylic layer was 92.4% at 450 nm (Fig. S2), as determined by spectroscopic measurement (UV-1800 UV-VIS, Shimadzu, Japan). The optical absorption of the PGS at 450 nm was 72.3% (Fig. S2). At maximum optical power, the power absorbed by the PGS was 1.63 W, accounting for all of the aforementioned losses.
Imaging
The PGS photoabsorber material and geometry were characterized using a microscope (AmScope, USA) and a laboratory scale with $50$ $\mu$m divisions. We used a MWIR camera (Fast M3K, Telops, USA) and microscope lens (1x, Telops, USA) to collect the thermal data in Fig. 2. The camera was directed at the sample via a mid-IR gold mirror (PFSQ20-03-M02, Thorlabs, USA) (Fig. S14A, B). The top elastomer sheet of the display was removed during thermal imaging. Time-resolved infrared videos were captured at a frame rate of 5,000 fps using an exposure time of 5.04 $\mu$s, and subsequently analyzed in RevealIR (Telops, USA) where an emissivity of 0.39 was used (Fig. S3). Radiometric temperature calibration was performed separately for low and high-temperature ranges, $T<236^{\circ}$ and $T>236^{\circ}$. Data captured across both ranges was combined in software (Fig. 2D), and aligned via temporal synchronization. For the sample with $w=0.55$ mm, we performed model-based calibration (Eq. (8) in Text S1) to infer temperatures $T>236^{\circ}$.
Video data was captured using a high-speed camera (Phantom VEO 640L, Vision Research, USA) with a microscope objective and a 2x Plan Apo Infinity Corrected lens (WE1603, Mitutoyo, USA). The camera was directed at the profile of each sample (Fig. S14C, D). A measured sample (thickness 2.2 mm) of cast acrylic was introduced into the field of view as a dimensional scaling reference. Optical imaging data were collected, reviewed, and saved in Phantom Video Player (Vision Research, USA). Metrology analysis was performed via manual annotation of points of interest using video analysis software (Tracker, Version 6.1.5), yielding displacement values (Fig. 2D-F).
An extended exposure photo, used for illustration in Figure 1, was captured using the camera of an iPhone SE 3rd generation via the app Spectre, with the ”light trail” feature, turned on. This image was not used for quantitative analysis.
Thermal resistance measurements
Values representing net effective thermal resistance $R$ to heat transfer in and out of the absorber, and absorber heat capacity $C$, were obtained from experimental data, by regression fit to the solution of a heat transfer equation (Text S2), with timescale $\tau=RC$. The fit was performed using temperature data acquired during the first 30 ms of photostimulation at 1.63 W. The temporal dependence of temperature, through $\tau=RC$, and the value of $R$ obtained from the fit, is generic and insensitive to modeling choices. The absorber was initially at room temperature during data collection. The experimentally determined heat capacity was $C=101\pm 16$ $\mu$ JK-1, which was in agreement with the value we calculated based on the geometry and properties of the photoabsorber, using $C=V_{abs}\rho_{abs}c_{abs}=85$ $\mu$ JK-1, where $V_{abs}$, $\rho_{abs}$, and $c_{abs}$ are the volume, density, and specific heat capacity of PGS optical target, respectively.
Numerical experiments
We studied the thermal and mechanical response characteristics of our system using finite element numerical analysis (FEA, COMSOL Multiphysics). The simulations accounted for heat transfer, gas expansion, and membrane mechanics. Detailed descriptions of the experiments and simulation methods are documented in Text S2.
Mechanical characterizations
Force measurements were captured with a load cell (LSB200 FSH02602, Futek, USA) and a strain gauge amplifier (CSG110, Futek, USA), using a data acquisition unit. The sample rate was 10 kHz. The load cell was applied with a preload of 0.74 N to ensure stable measurement conditions. Surface displacement measurements, as reported in Fig. 3, were captured via Laser Interferometer (Mikrotrak II LTC-025-02, MTI Instruments, USA) at a sample rate of 100 kHz.
Perceptual experiments
All six perceptual experiments were performed using a 437-pixel display surface (7.5 cm x 7.5 cm), with a pixel pitch of 3.6 mm (pixel density: 7.7 cm-2).
To inhibit light leakage, the flexible membrane (EcoFlex 00-10) was dyed white (Silc Pig White, Smooth-On), and a sheet of gold leaf ($\sim$0.1 $\mu$m thick) was adhered to the base. Neither modification significantly affected the membrane mechanical response. A black curtain occluded the display from participants’ view during the experiment. While we observed no stray light, participants wore laser safety glasses to avoid safety concerns.
All participants gave their written informed consent. The protocol was approved by the human subjects review board at the authors’ institution. The same ten participants (4 female, 6 male, age 19 to 34 years) completed the sequen ce of six experiments without training. None reported diagnosed health issues or ailments affecting touch sensation. One additional participant was unable to complete the perceptual tasks for personal reasons unrelated to the experiment or device. This data was excluded. Participants’ hand dimensions were measured. Right index finger pad widths were between 1.4 to 2.0 cm (mean: 1.6 cm), and hand lengths were between 16.5 to 19.5 cm (mean: 18.3 cm). Participants wore noise isolating circumaural headphones playing white noise to mask auditory cues during the experiment. For each experiment, participants finger were positioned as depicted in Figure 4. Before each experiment, participants were presented with each stimulus once for familiarization. Detailed descriptions of the design, method, protocol, and analyses used for each of the six perceptual experiments are reported in Text S5.
@@ Acknowledgments
We gratefully acknowledge Yangying Zhu, Shantal Adajian, Bolin Liao, and Alban Sauret for assistance with characterization experiments. We thank Neeli Tummala for valuable comments. We acknowledge the Innovation Workshop and Microfluidics Lab in the UCSB California NanoSystems Institute, the UCSB MRSEC facility (NSF DMR 0520415), and the UCSB Nanofabrication facility, which were used in this research.
Author Contributions: M.L. and Y.V. designed research; M.L, G.R., D.G. and Y.V. performed research; M.L. and V.K. analyzed data; M.L., E.H., and Y.V. wrote the paper.